# Thresholded Lasso for high dimensional variable selection

### v2

@@ 4.4 Variable selection in $A_{0}$
Note that $\beta_{\min,A_{0}}=\left\lVert\beta^{(11)}\right\rVert_{\infty}=C_{a}\lambda\sigma$, $\left\lVert\beta^{(12)}\right\rVert_{\infty}=C_{m}\lambda\sigma$, and $\left\lVert\beta^{(2)}\right\rVert_{\infty}=C_{t}\lambda\sigma=c_{t}\sigma/\sqrt{n}$. We mention up front that no matter where we put the threshold, some of the signals in $\beta^{(12)}$ will be lost so long as $t_{0}\asymp\lambda\sigma$ since $C_{m}\leq 1$; cf. Fig. 3. In our experiments, we can consistently recover those signals in $A_{0}$ for $\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|\asymp\lambda\sigma\sqrt{s_{0}}$ (in case $C_{a}=8.528$ for the model class in Figure 1), but this is not the case when $\beta_{\min,A_{0}}\asymp\lambda\sigma$ (in case $C_{a}=1.706$). This difference occurs despite the nearly identical $\ell_{p},p=1,2$ norm bounds on the estimation error for the initial Lasso estimator $\beta_{\text{\rm init}}$ for the two models of $\beta$; See the top two panels in Figure 2, where curves corresponding to different values of $\beta_{\min,A_{0}}$ with $C_{a}\lambda\sigma=0.2,1$ align well across different values of $s\in\{130,370,511\}$; This is true for $\left\lVert h_{T_{0}}\right\rVert_{2}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, and $\left\lVert\delta\right\rVert_{2}$ under the same design matrix $X$ of dimension 1600×20481600$\times$ 20481600 × 2048. The relative effect on variable selection in component $\beta^{(12)},\beta^{(2)}$, and $\beta^{(0)}$ follows the same trend for both $C_{a}$ settings. The relative effect on variable selection in component $\beta^{(11)}$ is much more significant in case $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$), since the largest error magnitude may reach the full signal strength, while for $\beta_{\min,A_{0}}=1$ ($C_{a}=8.528$), the magnitude of the error is only a fraction $\propto 1/{\sqrt{s_{0}}}$ of the signal strength in $A_{0}$ and hence $A_{0}\subset I$ so long as $t_{0}=o(\lambda\sigma\sqrt{s_{0}})$.
Table 1: Evaluation metrics for variable selection. $I$ is the estimated model, and ${\mathcal{D}}=I^{c}$.
@@ 5 Numerical results
[…]
@@
    Feed $Y$ and $X$ to the Thresholded Lasso algorithm to estimate $\beta$ as described in Section 1. We call the Lasso procedure $\textsf{LARS}(Y,X)$  [15] to compute the full regularization path. We select the $\beta_{\text{\rm init}}$ from this output path with penalty parameter $\lambda_{n}=f_{p}\lambda\sigma$. We then threshold $\beta_{\text{\rm init}}$ with threshold $t_{0}=f_{t}\lambda\sigma$, and run OLS to obtain $\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}$.
@@
Figure 2: $p=2048,n=1600$. Left column: $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, and $\left\lVert h\right\rVert_{1}$ as Lasso penalty ($f_{p}$) increases across different sparsity $s\in\{130,370,511\}$. Right column: plots of $\left\lVert h_{T_{0}}\right\rVert_{2}$ and $\left\lVert\delta\right\rVert_{2}$. In the top panel, we fix $\gamma=0.3$, and compare two cases of $C_{a}\lambda\sigma\in\{0.2,1\}$. In the middle panel, we fix $C_{a}\lambda\sigma=1$ and compare two cases of $\gamma\in\{0.3,0.7\}$. In the bottom panel, we zoom in on one case with $\gamma=0.7,C_{a}\lambda\sigma=0.2$, and we plot $\left\lVert\delta\right\rVert_{1}$ together with $\left\lVert\delta\right\rVert_{2}$ in the bottom right figure.
@@
[…]
@@ 5.1 $\ell_{1}$ and $\ell_{2}$ error bounds for $\beta_{\text{\rm init}}$
Dependence on $\gamma$. In the top row, all curves are closely aligned, validating that $\beta_{\min,A_{0}}$ (for different $C_{a}$s) does not impact these error metrics significantly. In contrast, in the middle row, we observe that curves for $\left\lVert h_{T_{0}}\right\rVert_{1}$ and $\left\lVert h_{T_{0}}\right\rVert_{2}$, shift downwards slightly for $\gamma=0.3$. This is expected since for the design matrix $X$ with smaller $\gamma$, the incoherence parameters appearing in Theorem 2.4 will be smaller. The two sets of curves for $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ and $\left\lVert h\right\rVert_{1}$ cross each other for $\gamma=0.3$ and 0.70.70.7 with a small gap (middle left panel), but eventually the set with a larger $\gamma$ stays on the top. The gap is potentially caused by the non-linear interactions between $\gamma$ and the penalty parameters throughout the entire Lasso path.
$\ell_{1}$ and $\ell_{2}$ error for $\delta$. In the right column, for the $\ell_{2}$ error for estimating $\beta$ with $\beta_{\text{\rm init}}$, we observe the typical V-shaped curves as $f_{p}$ increases from 0 to 111, since $\left\lVert\delta\right\rVert_{2}$ reaches a minimum and then increases again as the penalty $\lambda_{n}$ increases. In the bottom right panel of Fig. 2, we plot $\left\lVert\delta\right\rVert_{1}$ and $\left\lVert\delta\right\rVert_{2}$, and also $\left\lVert h\right\rVert_{1}$ (middle solid curves) for reference. All solid curves in the bottom corresponding to the same pair of $(s_{0},\gamma)$ again align well as predicted by Theorem 2.4 (bottom right panel). In the same panel, we observe the three dashed curves corresponding to $\left\lVert\delta\right\rVert_{1}$ are clearly separated under different sparsity $s\in\{130,370,511\}$, and stacked in descending order as $s$ decreases, that is, when $\beta$ becomes more sparse in (89). We know $\left\lVert\delta\right\rVert_{1}=\left\lVert\delta_{T_{0}}\right\rVert_{1}+\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$, where $\left\lVert\delta_{T_{0}}\right\rVert_{1}=\left\lVert h_{T_{0}}\right\rVert_{1}$. In contrast, as shown in the left panels, $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ and hence $\left\lVert h\right\rVert_{1}$ all align well across different $s$ for the same $\gamma$. Hence, the difference in $\left\lVert\delta\right\rVert_{1}$ is due to the component $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$, which depends on the sparsity. This phenomenon is expected and explained in Section 4.3. The influence of $\gamma$ on $\left\lVert\delta\right\rVert_{2}$ follows a similar trend as that for $\left\lVert h_{T_{0}}\right\rVert_{2}$. Hence, all error curves corresponding to a larger $\gamma$ (0.70.70.7) consistently dominate those with a smaller $\gamma=0.3$ for $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{2}$, and $\left\lVert\delta\right\rVert_{2}$ in the middle panels.
@@ 5.2 Variable selection with thresholding
In Figure 3, we plot in all panels the final model size $\left\lvert I\right\rvert$ (top right solid black curves), the number of TPs in model $I$ ($\left|I\cap T_{0}\right|$, middle diagonal red dashed lines/curves), and the number of FPs from $T_{0}^{c}$, ($\left|I\cap T_{0}^{c}\right|$, left bottom dotted green curves), as functions of threshold $t_{0}$ for $\beta_{\min,A_{0}}=C_{a}\lambda\sigma\in\{0.2,1\}$ and $C_{a}\in\{1.706,8.528\}$. In panel (a), all curves align well across different $s\in\{130,370,511\}$. At $t_{0}=C_{m}\lambda\sigma$ (middle dotted vertical line), the model size is only slightly above $\left\lvert A_{0}\right\rvert=30$ for both cases of $C_{a}$. When the top curves (plotting $\left|I\right|$) touch upon or cross over the horizontal line of $y=30$ at $t_{0}=t^{\prime}$, where $t^{\prime}\in(\lambda\sigma/2,\lambda\sigma)$, the model contains $A_{0}$ exactly. For $\beta_{\min,A_{0}}=1$, all coordinates in $\beta^{(11)}$ remain in model $I$ so long as $t_{0}\leq 2\lambda\sigma$. For $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$), the model size $\left\lvert I\right\rvert$ continues to shrink till 0, as $t_{0}$ increases.
[…]
@@
False negatives. Recall FNs = $s_{0}$ TPs. The primary distinction between the two settings of $C_{a}$ lies in the FNs from $\beta^{(11)}$, since we will lose some variables from $\beta^{(12)}$ inevitably for both choices of $C_{a}$, no matter where we put $t_{0}$, as $\beta^{(12)}_{\text{\rm init},j}$ may fall within the range of ±$\pm c\lambda\sigma$ for any $c\in(0,1]$. The entries of the Lasso estimate $\beta_{\text{\rm init}}^{(12)}$ of $\beta^{(12)}$ are indeed spread across the interval of $[-\lambda\sigma,\lambda\sigma]$ since $\left\lvert\beta_{\text{\rm init},j}^{(12)}\right\rvert\leq\lambda\sigma$. This is indicated in Fig. 3 by the negative slope in the red diagonal lines, where coordinates in $\beta^{(12)}$ are regularly cut as $t_{0}$ increases. Larger $\beta_{\min,A_{0}}$ means variables in $A_{0}$ will be kept over a longer range of $t_{0}$, as the remaining TPs are all from $A_{0}$ after a certain threshold. Indeed, for $\beta_{\min,A_{0}}=1$ ($C_{a}=8.528$), we observe in Fig. 3 panel (c), the curves for TPs (red dashed slanted lines) have a changing slope at around $t_{0}=t^{\prime}$, where $t^{\prime}\in(\lambda\sigma/2,\lambda\sigma]$ for $f_{p}=0.3$, and then flattens out along the horizontal line of $y=30$ until $t_{0}=4\lambda\sigma$. For $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$, panel (d)), the dashed diagonal line intersects the horizontal line of $y=30$, and continues with the downward trend until it reaches $y=0$ while losing all true variables.
@@
False positives. We observe in Fig. 3 panel (a) that FPs drop sharply in both $C_{a}$ cases, and the rate is the same for all $s\in\{130,511,770\}$ as $t_{0}$ increases, whereas TPs drop with a slower rate due to their larger estimated values. At $t_{0}=\lambda\sigma/2$, the model size is about but slightly below 505050 with both FNs and FPs. By our theory, the coordinates in $\beta^{(2)}$ with small coefficients are at the noise level and hence are neither guaranteed nor necessary to be included in the model $I$. Roughly speaking, the largest magnitude of the Lasso estimate $\beta_{\text{\rm init}}^{(2)}$ and $\beta_{\text{\rm init}}^{(0)}$ as well as their $\ell_{\infty}$ norm error are nearly all bounded by $\lambda\sigma$ in absolute values, and hence $t_{0}\asymp\lambda\sigma$ is effective in controlling the number of variables selected from $T_{0}^{c}$ (False Positives, cf. Table 1), as shown on the left bottom dotted green curves on all panels in Figure 3. On each curve, the number of FPs drops quickly as $t_{0}$ increases and transitions to a horizontal line at $t_{0}\approx 0.8\lambda\sigma$, as predicted by the lower bound on $t_{0}$ in (69).
Refer to caption Refer to caption
Figure 4: $p=2048,n=1600$, $s=130$. Plots of $\left\lVert\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.54248pt}}-\beta\right\rVert_{2}$, for $C_{a}\lambda\sigma\in\{0.2,1\}$ and $\gamma\in\{0.3,0.7\}$. The horizontal lines correspond to the $\ell_{2}$-norm error of Lasso estimate $\beta_{\text{\rm init}}$, namely, $\left\lVert\delta\right\rVert_{2}$.
5.3 $\ell_{2}$-norm error of $\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}$
@@
In this section, we show that the previously stated tradeoffs between FPs and TPs in Thresholded Lasso does not come at the cost of an increased $\ell_{2}$ error of the final estimator. We plot in Fig. 4, the $\ell_{2}$-norm error of the final estimate, defined as $\left\lVert\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}-\beta\right\rVert_{2}$, we observe that for all the cases of Lasso penalty ($f_{p}$ values), for a wide range of $t_{0}$, $\left\lVert\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}-\beta\right\rVert_{2}$ stays at the same level or below $\left\lVert\delta\right\rVert_{2}$. This is as predicted by Theorems 2.1 and 2.4. For lower $f_{p}$ values, such as 0.1,0.30.1,0.30.1 , 0.3, $\left\lVert\hat{\beta}_{I}^{\mathop{\text{\rm ols}\kern 0.60275pt}}-\beta\right\rVert_{2}$ further decreases as threshold $t_{0}$ increases. This is because small coordinates in $\beta_{\text{\rm init}}$ with values below $t_{0}$ will be gone with thresholding, while variables in $A_{0}$ remain intact due to their larger magnitudes. When $f_{p}$ increases from 0.30.30.3 to 111, $\left\lVert\delta\right\rVert_{2}$ has a V-shaped curve, due to the loss of significant coordinates in $T_{0}$. This is shown in Fig. 4, where the horizontal lines corresponding to $f_{p}=1$ is higher than the ones correspond to $f_{p}=0.5$ or 0.70.70.7; See also Fig. 2. In practice, we recommend using cross-validation to set the Lasso penalty $\lambda_{n}$, which typically ranges between $0.3\lambda\sigma$ and $0.7\lambda\sigma$ for this example, and then apply thresholding.
6 Conclusion
Table 2: $\beta$ configurations with fixed $a_{0}=30,s_{0}=50.$
@@ 6 Conclusion
Table 3: In this table, we list the actual values of $s$ generated through (79). We also list magnitudes of each $\beta$ component and their $\ell_{1}$ and $\ell_{2}$ norms. Here $c_{t}=C_{t}\sqrt{2\log p}$. The component $\beta^{(11)}=\beta_{A_{0}}$ has $a_{0}=30$ non-zero coordinates with magnitude $C_{a}\lambda\sigma$, where $C_{a}>1$. The component $\beta^{(12)}=\beta_{T_{0}\setminus A_{0}}$ has $s_{0}-a_{0}=20$ non-zero coordinates with magnitude $C_{m}\lambda\sigma$, where $C_{m}=1/{\sqrt{2}}$ for $s>s_{0}$. The component $\beta^{(2)}=\beta_{S\setminus T_{0}}$ consists of $s-s_{0}$ non-zero coordinates of magnitude $c_{t}\sigma/\sqrt{n}$, where $c_{t}$ is an absolute constant.
Acknowledgement
I dedicate this work to Professor Peter Bühlmann on the occasion of his 60th birthday. His unwavering support and mentorship during my career has been deeply appreciated. I thank Peter Bühlmann, Emmanuel Candès, Sara van de Geer, John Lafferty, Po-Ling Loh, Richard Samworth, Xiaotong Shen, Martin Wainwright, Larry Wasserman, Dana Yang and Cun-Hui Zhang for helpful discussions and encouragement. The author thanks Bin Yu warmly for hosting her visit at UC Berkeley while she finishes this work in Spring 2010. Part of this work has appeared in a conference paper with title: Thresholding Procedures for High Dimensional Variable Selection and Statistical Estimation, in Proceedings of Advances in Neural Information Processing Systems 22.
Appendix A Proofs
A.1 Proof of Theorem 2.1

### v3

@@ 4.4 Variable selection in $A_{0}$
Note that $\beta_{\min,A_{0}}=\left\lVert\beta^{(11)}\right\rVert_{\infty}=C_{a}\lambda\sigma$, $\left\lVert\beta^{(12)}\right\rVert_{\infty}=C_{m}\lambda\sigma$, and $\left\lVert\beta^{(2)}\right\rVert_{\infty}=C_{t}\lambda\sigma=c_{t}\sigma/\sqrt{n}$. We mention up front that no matter where we put the threshold, some of the signals in $\beta^{(12)}$ will be lost so long as $t_{0}\asymp\lambda\sigma$ since $C_{m}\leq 1$; cf. Fig. 3. In our experiments, we can consistently recover those signals in $A_{0}$ for $\beta_{\min,A_{0}}:=\min_{j\in A_{0}}|\beta_{j}|\asymp\lambda\sigma\sqrt{s_{0}}$ (in case $C_{a}=8.528$ for the model class in Figure 1), but this is not the case when $\beta_{\min,A_{0}}\asymp\lambda\sigma$ (in case $C_{a}=1.706$). This difference occurs despite the nearly identical $\ell_{p},p=1,2$ norm bounds on the estimation error for the initial Lasso estimator $\beta_{\text{\rm init}}$ for the two models of $\beta$; See the top two panels in Figure 2, where curves corresponding to different values of $\beta_{\min,A_{0}}$ with $C_{a}\lambda\sigma=0.2,1$ align well across different values of $s\in\{130,370,511\}$; This is true for $\left\lVert h_{T_{0}}\right\rVert_{2}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, and $\left\lVert\delta\right\rVert_{2}$ under the same design matrix $X$ of dimension $1600\times 2048$. The relative effect on variable selection in component $\beta^{(12)},\beta^{(2)}$, and $\beta^{(0)}$ follows the same trend for both $C_{a}$ settings. The relative effect on variable selection in component $\beta^{(11)}$ is much more significant in case $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$), since the largest error magnitude may reach the full signal strength, while for $\beta_{\min,A_{0}}=1$ ($C_{a}=8.528$), the magnitude of the error is only a fraction $\propto 1/{\sqrt{s_{0}}}$ of the signal strength in $A_{0}$ and hence $A_{0}\subset I$ so long as $t_{0}=o(\lambda\sigma\sqrt{s_{0}})$.
Table 1: Evaluation metrics for variable selection. II is the estimated model, and ${\mathcal{D}}=I^{c}$.
@@ 5 Numerical results
In this section, we present results from numerical simulations to validate the theoretical analysis presented in previous sections. We consider Gaussian random matrices for the design $X$ with both $p\times p$ identity and Toeplitz covariance. We refer to the former as i.i.d. Gaussian ensemble, where $X_{ij}\sim N(0,1)$ for all $i,j$, and the latter as Toeplitz ensemble, where the covariance matrix $T(\gamma)$ for each row vector in $\mathbb{R}^{p}$ is given by $T(\gamma)]_{i,j}=\gamma^{|i-j|}$. The design matrix dimensions are either $(p=1024,n=800)$ or $(p=2048,n=1600)$. To evaluate the impact of nominal sparsity $s$ on the recovery of $s_{0}$ components, we use $\beta$ as constructed as in Section 4.2. We have the following steps.
@@
    Feed $Y$ and $X$ to the Thresholded Lasso algorithm to estimate $\beta$ as described in Section 1. We call the Lasso procedure $\textsf{LARS}(Y,X)$  [15] to compute the full regularization path. We select the $\beta_{\text{\rm init}}$ from this output path with penalty parameter $\lambda_{n}=f_{p}\lambda\sigma$. We then threshold $\beta_{\text{\rm init}}$ with threshold $t_{0}=f_{t}\lambda\sigma$, and run OLS to obtain $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$.
@@
Figure 2: $p=2048,n=1600$. Left column: $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{1}$, and $\left\lVert h\right\rVert_{1}$ as Lasso penalty ($f_{p}$) increases across different sparsity $s\in\{130,370,511\}$. Right column: plots of $\left\lVert h_{T_{0}}\right\rVert_{2}$ and $\left\lVert\delta\right\rVert_{2}$. In the top panel, we fix $\gamma=0.3$, and compare two cases of $C_{a}\lambda\sigma\in\{0.2,1\}$. In the middle panel, we fix $C_{a}\lambda\sigma=1$ and compare two cases of $\gamma\in\{0.3,0.7\}$. In the bottom panel, we zoom in on one case with $\gamma=0.7,C_{a}\lambda\sigma=0.2$, and we plot $\left\lVert\delta\right\rVert_{1}$ together with $\left\lVert\delta\right\rVert_{2}$ in the bottom right panel.
@@
Figure 3: $p=2048,n=1600,\gamma=0.7$. Plots of model size ($|I|$), number of TPs and FPs, as threshold increases. Note $|I|=$ TPs + FPs. In (a) and (b), Lasso penalty factor $f_{p}=0.3$ is fixed, and in panel (a) $s\in\{130,511,710\}$, and in panel (b) $s\in\{50,130\}$. In panels (c) and (d), we plot the same metrics across different $f_{p}\in\{0.1,0.3,0.7\}$ with fixed $s=130$. In all panels, the 3 dotted vertical lines from left to right represent $C_{m}\lambda\sigma/2,C_{m}\lambda\sigma$ and $\lambda\sigma$. The model size remains invariant and hence the diagonal dashed lines all stay flat for $\lambda\sigma<t_{0}\leq 2\lambda\sigma$ for $\beta_{\min,A_{0}}=1$.
@@ 5.1 $\ell_{1}$ and $\ell_{2}$ error bounds for $\beta_{\text{\rm init}}$
Dependence on $\gamma$. In the top row, all curves are closely aligned, validating that $\beta_{\min,A_{0}}$ (for different values of $C_{a}$) does not impact these error metrics significantly. In contrast, in the middle row, we observe that curves for $\left\lVert h_{T_{0}}\right\rVert_{1}$ and $\left\lVert h_{T_{0}}\right\rVert_{2}$, shift downwards slightly for $\gamma=0.3$. This is expected since for the design matrix $X$ with smaller $\gamma$, the incoherence parameters appearing in Theorem 2.4 will be smaller. The two sets of curves for $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ and $\left\lVert h\right\rVert_{1}$ cross each other for $\gamma=0.3$ and $0.7$ with a small gap (middle left panel), but eventually the set with a larger $\gamma$ stays on the top. The gap is potentially caused by the non-linear interactions between $\gamma$ and the penalty parameters throughout the entire Lasso path.
$\ell_{1}$ and $\ell_{2}$ error for $\delta$. In the right column, for the $\ell_{2}$ error for estimating $\beta$ with $\beta_{\text{\rm init}}$, we observe the typical V-shaped curves as $f_{p}$ increases from 0 to $1$, since $\left\lVert\delta\right\rVert_{2}$ reaches a minimum and then increases again as the penalty $\lambda_{n}$ increases. In the bottom right panel of Fig. 2, we plot $\left\lVert\delta\right\rVert_{1}$ and $\left\lVert\delta\right\rVert_{2}$, and also $\left\lVert h\right\rVert_{1}$ (middle solid curves) for reference. All solid curves in the bottom corresponding to the same pair of $(s_{0},\gamma)$ again align well as predicted by Theorem 2.4 (bottom right panel). In the same panel, we observe the three dashed curves corresponding to $\left\lVert\delta\right\rVert_{1}$ are clearly separated under different sparsity $s\in\{130,370,511\}$, and stacked in descending order as $s$ decreases, that is, when $\beta$ becomes more sparse in (89). We know $\left\lVert\delta\right\rVert_{1}=\left\lVert\delta_{T_{0}}\right\rVert_{1}+\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$, where $\left\lVert\delta_{T_{0}}\right\rVert_{1}=\left\lVert h_{T_{0}}\right\rVert_{1}$. In contrast, as shown in the left panels, $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}^{c}}\right\rVert_{1}$ and hence $\left\lVert h\right\rVert_{1}$ all align well across different $s$ for the same $\gamma$. Hence, the difference in $\left\lVert\delta\right\rVert_{1}$ is due to the component $\left\lVert\delta_{T_{0}^{c}}\right\rVert_{1}$, which depends on the sparsity. This phenomenon is expected and explained in Section 4.3. The influence of $\gamma$ on $\left\lVert\delta\right\rVert_{2}$ follows a similar trend as that for $\left\lVert h_{T_{0}}\right\rVert_{2}$. Hence, all error curves corresponding to a larger $\gamma$ ($0.7$) consistently dominate those with a smaller $\gamma=0.3$ for $\left\lVert h_{T_{0}}\right\rVert_{1}$, $\left\lVert h_{T_{0}}\right\rVert_{2}$, and $\left\lVert\delta\right\rVert_{2}$ in the middle panel.
@@ 5.2 Variable selection with thresholding
In Figure 3, we plot in all panels the final model size $\left\lvert I\right\rvert$ (top right solid black curves), the number of TPs in model II ($\left|I\cap T_{0}\right|$, middle diagonal red dashed lines/curves), and the number of FPs from $T_{0}^{c}$, ($\left|I\cap T_{0}^{c}\right|$, left bottom dotted green curves), as functions of threshold $t_{0}$ for $\beta_{\min,A_{0}}=C_{a}\lambda\sigma\in\{0.2,1\}$ and $C_{a}\in\{1.706,8.528\}$. In panel (a), all curves align well across different $s\in\{130,370,511\}$. At $t_{0}=C_{m}\lambda\sigma$ (middle dotted vertical line), the model size is only slightly above $\left\lvert A_{0}\right\rvert=30$ for both cases of $C_{a}$. When the top curves (plotting $\left|I\right|$) touch upon or cross over the horizontal line of $y=30$ at $t_{0}=t^{\prime}$, where $t^{\prime}\in(\lambda\sigma/2,\lambda\sigma)$, the model contains $A_{0}$ exactly. For $\beta_{\min,A_{0}}=1$, all coordinates in $\beta^{(11)}$ remain in model $I$ so long as $t_{0}\leq 2\lambda\sigma$. For $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$), the model size $\left\lvert I\right\rvert$ continues to shrink till 0, as $t_{0}$ increases.
In Fig.  3 panel (b), we compare the exact $s_{0}$-sparse case ($s=50$) with the almost $s_{0}$-sparse case ($s=130$). Under the same $t_{0}$, we observe that the exact sparse case recovers more non-zero components from $T_{0}$ (higher red dashed lines with more TPs and fewer FNs for $t_{0}<1.5\lambda\sigma$) and less from $T_{0}^{c}$ (lower green lines with fewer FPs for $t_{0}<\lambda\sigma$), since for $s=s_{0}$, $\beta_{j}^{(12)}$ is larger in magnitude with $C_{m}=1$ and $\beta_{T_{0}^{c}}=0$. Panels (c) and (d) show that as the Lasso penalty ($f_{p}$) increases, model sizes further decrease, and hence $\left\lvert I\cap T_{0}\right\rvert$ and $\left\lvert I\cap T_{0}^{c}\right\rvert$ both decrease. Hence all three sets of curves ($|I|$, TPs, FPs) shift downward as $f_{p}$ increases since Lasso is able to remove some less significant variables as an initial estimator. However, FPs remain at a high level without thresholding or when the threshold is small. This is true for both $C_{a}$ settings.
@@
False negatives. Recall FNs = $s_{0}$ - TPs. The primary distinction between the two settings of $C_{a}$ lies in the FNs from $\beta^{(11)}$, since we will lose some variables from $\beta^{(12)}$ inevitably for both choices of $C_{a}$, no matter where we put $t_{0}$, as $\beta^{(12)}_{\text{\rm init},j}$ may fall within the range of $\pm c\lambda\sigma$ for any $c\in(0,1]$. The entries of the Lasso estimate $\beta_{\text{\rm init}}^{(12)}$ of $\beta^{(12)}$ are indeed spread across the interval of $\lambda\sigma,\lambda\sigma]$ since $\left\lvert\beta_{\text{\rm init},j}^{(12)}\right\rvert\leq\lambda\sigma$. This is indicated in Fig. 3 by the negative slope in the red diagonal lines, where coordinates in $\beta^{(12)}$ are regularly cut as $t_{0}$ increases. Larger $\beta_{\min,A_{0}}$ means variables in $A_{0}$ will be kept over a longer range of $t_{0}$, as the remaining TPs are all from $A_{0}$ after a certain threshold. Indeed, for $\beta_{\min,A_{0}}=1$ ($C_{a}=8.528$), we observe in Fig. 3 panel (c), the curves for TPs (red dashed slanted lines) have a changing slope at around $t_{0}=t^{\prime}$, where $t^{\prime}\in(\lambda\sigma/2,\lambda\sigma]$ for $f_{p}=0.3$, and then flattens out along the horizontal line of $y=30$ until $t_{0}=4\lambda\sigma$. For $\beta_{\min,A_{0}}=0.2$ ($C_{a}=1.706$, panel (d)), the dashed diagonal line intersects the horizontal line of $y=30$, and continues with the downward trend until it reaches $y=0$ while losing all true variables.
@@
False positives. We observe in Fig. 3 panel (a) that FPs drop sharply in both $C_{a}$ cases, and the rate is the same for all $s\in\{130,511,770\}$ as $t_{0}$ increases, whereas TPs drop with a slower rate due to their larger estimated values. At $t_{0}=\lambda\sigma/2$, the model size is about but slightly below $50$ with both FNs and FPs. By our theory, the coordinates in $\beta^{(2)}$ with small coefficients are at the noise level and hence are neither guaranteed nor necessary to be included in the model $I$. Roughly speaking, the largest magnitude of the Lasso estimate $\beta_{\text{\rm init}}^{(2)}$ and $\beta_{\text{\rm init}}^{(0)}$ as well as their $\ell_{\infty}$ norm error are nearly all bounded by $\lambda\sigma$ in absolute values, and hence $t_{0}\asymp\lambda\sigma$ is effective in controlling the number of variables selected from $T_{0}^{c}$ (False Positives, cf. Table 1), as shown on the left bottom dotted green curves on all panels in Figure 3. On each curve, the number of FPs drops quickly as $t_{0}$ increases and transitions to a horizontal line at $t_{0}\approx 0.8\lambda\sigma$, as predicted by the lower bound on $t_{0}$ in (66).
@@
$\ell_{2}$-norm error of $\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)$. In this section, we show that the previously stated tradeoffs between FPs and TPs in Thresholded Lasso does not come at the cost of an increased $\ell_{2}$ error of the final estimator. We plot in Fig. 4, the $\ell_{2}$-norm error of the final estimate, defined as $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}$, we observe that for all the cases of Lasso penalty ($f_{p}$ values), for a wide range of $t_{0}$, $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}$ stays at the same level or below $\left\lVert\delta\right\rVert_{2}$. This is as predicted by Theorems 2.1 and 2.4. For lower $f_{p}$ values, such as $0.1,0.3$, $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.60275pt}}(I)-\beta\right\rVert_{2}$ further decreases as threshold $t_{0}$ increases. This is because small coordinates in $\beta_{\text{\rm init}}$ with values below $t_{0}$ will be gone with thresholding, while variables in $A_{0}$ remain intact due to their larger magnitudes. When $f_{p}$ increases from $0.3$ to $1$, $\left\lVert\delta\right\rVert_{2}$ has a V-shaped curve, due to the loss of significant coordinates in $T_{0}$. This is shown in Fig. 4, where the horizontal lines corresponding to $f_{p}=1$ is higher than the ones correspond to $f_{p}=0.5$ or $0.7$; See also Fig. 2.
@@
Table 2: In this table, we list the actual values of $s$ generated through (76) for $\beta$ configurations with $a_{0}=30,s_{0}=50.$ We also list magnitudes of each $\beta$ component and their $\ell_{1}$ and $\ell_{2}$ norms. Here $c_{t}=C_{t}\sqrt{2\log p}$. The component $\beta^{(11)}=\beta_{A_{0}}$ has $a_{0}=30$ non-zero coordinates with magnitude $C_{a}\lambda\sigma$, where $C_{a}>1$. The component $\beta^{(12)}=\beta_{T_{0}\setminus A_{0}}$ has $s_{0}-a_{0}=20$ non-zero coordinates with magnitude $C_{m}\lambda\sigma$, where $C_{m}=1/{\sqrt{2}}$ for $s>s_{0}$. The component $\beta^{(2)}=\beta_{S\setminus T_{0}}$ consists of $s-s_{0}$ non-zero coordinates of magnitude $c_{t}\sigma/\sqrt{n}$, where $c_{t}$ is an absolute constant.
Refer to caption Refer to caption
Figure 4: $p=2048,n=1600$, $s=130$. Plots of $\left\lVert\hat{\beta}^{\mathop{\text{\rm ols}\kern 0.54248pt}}(I)-\beta\right\rVert_{2}$, for $C_{a}\lambda\sigma\in\{0.2,1\}$ and $\gamma\in\{0.3,0.7\}$. The horizontal lines correspond to the $\ell_{2}$-norm error of Lasso estimate $\beta_{\text{\rm init}}$, namely, $\left\lVert\delta\right\rVert_{2}$.
In practice, we recommend using cross-validation to set the Lasso penalty $\lambda_{n}$, which typically ranges between $0.3\lambda\sigma$ and $0.7\lambda\sigma$ for this example, and then apply thresholding.
6 Proofs
6.1 Proof of Theorem 2.1